one of the rectangle components of a velocity of 20ms-1 is 10ms-1  .find the other component.
 
 

Grade:11

1 Answers

Arun
5 years ago
given velocity v = 20 m/s

let the given rectangular component is

cos somponent, so 10 = v cos x

so, other component will be sin component

i.e., v sin x

or, 10 = 20 cos x

or, cos x = 1/2 

and as sin^2 x + cos^2 x = 1

or, 1 - cos^2 x = sin^2 x

or, 1 - (1/2)^2 = sin^2 x

so, sin x = sqrt( 1 - 1/4)

or, sin x = sqrt(3/4)

so, rectangular component is 

20 sin x = 20 sqrt(3/4)

= 10 sqrt3

= 17.32 m/s